### v2

# GPT4Scene: Understand 3D Scenes from Videos with Vision-Language Models

Abstract
In recent years, 2D Vision-Language Models (VLMs) have made significant strides in image-text understanding tasks. However, their performance in 3D spatial comprehension, which is critical for embodied intelligence, remains limited. Recent advances have leveraged 3D point clouds and multi-view images as inputs, yielding promising results. However, we propose exploring a purely vision-based solution inspired by human perception, which merely relies on visual cues for 3D spatial understanding. This paper empirically investigates the limitations of VLMs in 3D spatial knowledge, revealing that their primary shortcoming lies in the lack of global-local correspondence between the scene and individual frames. To address this, we introduce GPT4Scene, a novel visual prompting paradigm in VLM training and inference that helps build the global-local relationship, significantly improving the 3D spatial understanding of indoor scenes. Specifically, GPT4Scene constructs a 3D Bird’s Eye View (BEV) image from the video and marks consistent object IDs across both frames and the BEV image. The model then inputs the concatenated BEV image and video frames with markers. In zero-shot evaluations, GPT4Scene improves performance over closed-source VLMs like GPT-4o. Additionally, we prepare a processed video dataset consisting of 165K text annotation to fine-tune open-source VLMs, achieving state-of-the-art performance on all 3D understanding tasks. Surprisingly, after training with the GPT4Scene paradigm, VLMs consistently improve during inference, even without visual prompting and BEV image as explicit correspondence. It demonstrates that the proposed paradigm helps VLMs develop an intrinsic ability to understand 3D scenes, which paves the way for a noninvasive approach to extending pre-trained VLMs for 3D scene understanding.
Figure 1: The overall architecture of GPT4Scene. It is capable of understanding 3D scenes and performing tasks such as 3D question answering, dense captioning, and visual grounding using only video input. In contrast to 3D point LLMs, GPT4Scene takes input solely from the vision modality, with global information provided by the BEV image reconstructed from the 3D structure derived from the video.
1 Introduction
Embodied AI refers to intelligent systems capable of performing various tasks through interaction with the physical environment [41, 68, 10]. It has broad applications in industrial inspection, smart homes, and smart cities [41, 21, 10]. 3D scene understanding involves the ability of multi-modal language models to comprehend the overall layout of indoor environments and the spatial relationships between objects [6, 22, 13]. Thus, a strong foundation for embodied intelligence lies in understanding scenes effectively.
Currently, 3D point LLMs are a popular approach for understanding indoor scenes, using point clouds as input and aligning point features with LLMs to perform scene understanding tasks [21, 38]. However, this approach has several limitations. First, point clouds provide limited detailed information, such as fine geometric details, material properties, and intricate textures. Second, although some point LLMs [121, 21, 40] attempt to use point clouds and multi-images as inputs, they face challenges in aligning text, image, and point cloud modalities. Additionally, the significant imbalance in the data quantities of point clouds versus text/video data introduces further complications.
These limitations motivate us to explore indoor scene understanding using pure vision inputs. This approach aligns more closely with human perception, as people can understand 3D scenes without relying on explicit 3D data like point clouds. Vision Language Models (VLMs) have demonstrated impressive results in image-text multi-modal tasks[90, 62, 54, 119]. However, their application to understanding immersive 3D indoor scenes has not been well explored. We conduct a pilot study to investigate this potential by directly inputting a scene video into VLMs. Our results show that this approach causes VLMs to fail to understand 3D scenes. The core issue lies in the lack of global scene information and the misalignment between each frame’s local position and international context.
To address this, we propose a framework called GPT4Scene to help VLMs establish spatial relationships as shown in Figure 1. First, we perform 3D reconstruction based on the input video to generate a Bird’s Eye View (BEV) image as additional input, providing a comprehensive perspective of the scene layout. We also incorporate Spatial-Temporal Object markers (STO markers) in both the 3D BEV image and 2D frames. These STO markers represent object IDs consistent across sequential frames (temporal level) and align with those in the 3D BEV image (spatial level), helping the VLMs establish the global-local relationship between the overall scene layout and the video frames.
In a zero-shot setting, we first apply GPT4Scene to powerful closed-source VLMs, such as GPT-4o, using video frames and the BEV image with STO markers. The performance is comparable to existing 3D point-based LLMs, demonstrating GPT4Scene’s effectiveness in enhancing VLMs’ 3D scene understanding capabilities. For smaller open-source VLMs, we construct a ScanAlign dataset, which includes video frames with STO markers, BEV images, and text annotations. After fine-tuning on this dataset, our method outperforms existing approaches. Moreover, after fine-tuning, VLMs exhibit impressive performance with only the raw scene video as input, whereas additional processing was necessary for good results before training. It indicates that GPT4Scene helps VLMs develop an intrinsic ability to model 3D scene relationships.
Our paper makes these major contributions:
  • We introduce GPT4Scene, a framework that enhances Vision-Language Models (VLMs) to comprehend 3D scenes directly from pure vision input.
  • We introduce two techniques: (1) A 3D BEV image with global context information and (2) Spatial-Temporal Object markers (STO markers) for spatial and temporal consistency across BEV image and video frames.
  • We construct ScanAlign, a dataset consisting of video frames, BEV images with STO markers, and text annotations. Fine-tuning VLMs on this dataset markedly enhances their 3D scene understanding abilities.
  • GPT4Scene demonstrates robust performance in zero-shot and fine-tuning settings, achieving SOTA results across diverse 3D scene understanding tasks.
2 Related Work
3D indoor scene understanding. 3D indoor scene understanding allows robots to identify object positions, structures, and relationships within indoor environments, enabling question-and-answer interactions about the scene’s content. This process combines 3D perception with large language models (LLMs). 3D perception, as a foundational component, is typically trained on common indoor datasets [25, 8, 11, 69, 78, 89, 84, 28, 52, 116] using point clouds as input, supporting tasks like 3D object detection and instance segmentation [74, 48, 70, 82, 88, 96, 110]. Recent advancements in 3D Vision-Language Learning (3D-VL) combine 3D perception tasks with natural language, introducing diverse textual annotations on datasets like ScanNet to support tasks such as 3D Question Answering [6, 66, 106, 18], 3D Dense Captioning [22], and 3D Visual Grounding [13, 114, 2, 1, 51]. Initial studies focus on single 3D-VL tasks [108, 20, 97, 7, 36, 44, 46, 64, 104, 115, 33], while recent research introduc unified models for multiple tasks [9, 14]. The success of 2D-VL models is largely due to extensive 2D Vision-Language Pretraining (2D-VLP) with image-text pairs, leading some approaches to leverage 2D-VLP for 3D-VL [35, 37, 72, 86, 101, 112, 113, 99]. Though 3D-VLP has emerged [120, 121, 49, 29, 102, 47, 91], the limited availability of 3D point cloud-text pairs constrains its effectiveness. Therefore, achieving better results while utilizing 2D modality input would be beneficial.
3D Point Cloud LLMs. 3D Vision Language tasks aim to integrate 3D scene understanding with natural language processing while incorporating LLMs, which enables more natural human-computer interaction. 3D point LLMs take point clouds as input, supporting natural language generation and interaction in 3D scenes. Early 3D LLMs focused on object-level geometry and appearance [34, 76, 100, 75]. Later, 3D LLMs expanded to indoor scenes, emphasizing spatial relationships among objects and overall scene features, often using cene point clouds with auxiliary 2D multi-view images [38, 21, 94, 32, 68]. To better capture object relationships, recent 3D LLMs decouple scene objects before inputting them into LLMs [41, 40]. However, they still lack detailed information about object materials and precise geometric details. Additionally, they miss the current first-person perspective, whereas existing embodied AI can determine the current position within a room based on the current frame [10, 118, 60]. Moreover, due to the limited amount of point cloud data, its general understanding capabilities are relatively limited compared to current VLMs. Therefore, using the video modality to understand 3D scenes has more practical value.
Vision Language Models (VLMs). Vision Language Models (VLMs) are multimodal models that integrate visual and language processing capabilities, enabling the understanding and generation of combined image-text information. The origin of VLMs can be traced back to CLIP’s 2D image-text pair pretraining [77, 12, 81], which laid the foundation for incorporating LLMs. Early VLMs used attention mechanisms or Q-Former to fuse image and text modalities before inputting them to LLMs [55, 56, 27, 3, 45]. Later, an approach emerged that directly projects image features into the LLM using an MLP [61, 62, 54, 119, 90], achieving better performance. Building on this, VLMs expanded into visual grounding tasks [92, 73, 79, 53, 30, 85, 16] and further to video understanding by using spatiotemporal compression to process information from long image sequences [111, 24, 67, 59, 65, 4, 58, 83, 80, 5, 63]. However, VLMs cannot directly understand 3D scenes, so our task is to assist them in interpreting the 3D world.
3 Pilot Study
Figure 2: Pilot Study. We tested zero-shot 3D question answering on open-sourced VLM (Qwen2-VL) and closed-sourced VLM (GPT-4o). For GPT-4o, adding BEV images and corresponding object markers alongside video frames improved performance, but no significant improvement was seen with Qwen2-VL. While providing global features and connections between global and local information is helpful, smaller VLMs may need finetuning to enhance their 3D understanding capabilities. ‘VID’ means indoor video frames, ‘BEV’ means indoor BEV Map, and ‘Mrks’ means object ID markers on video frames and BEV maps.
Figure 3: The Framework of GPT4Scene. A scene video is processed by sampling frames, reconstructing a point cloud, and generating a BEV image. Object locations are detected from the point cloud and projected onto the video frames. The resulting frames and BEV image, enhanced with STO-Markers, are inputs for VLM training and inference.
In this paper, we opt for purely vision-based VLMs over 3D point LLMs. One reason is that 3D point clouds lack significant fine details. Another reason is that incorporating an additional 3D point cloud feature representation into pre-trained VLMs requires finetuning extensive point cloud data, which risks disrupting the well-established text-image alignment and potentially impairing the vision-language understanding capability. Additionally, accepting pure first-person perspective video input aligns more closely with the nature of human perception. Therefore, we choose VLMs to understand 3D scenes. The next question is: can VLMs effectively understand the 3D world directly?
Can VLMs understand the 3D world directly? To explore this issue, we conducted a pilot study, and the results are shown in Figure 2. We input a preprocessed scene video into the VLMs for the 3D question answering task. We find that neither the open-source Qwen2-VL-7B nor the powerful closed-source GPT-4o could perform the task effectively. We find that VLMs perform well in answering specific details about the scene but struggle with understanding the spatial relationships between objects. Therefore, we aim to provide VLMs with global information. We add the BEV image rendered from the indoor scene reconstructed from the input video. We find that this approach led to significant improvements in GPT-4o’s performance. Inspired by SOM [103], which enhances VLMs’ understanding by marking objects on images, we further explore whether providing markers on the video frames could improve VLMs’ spatial learning. We label the objects in the scene, marking them on both the video frames and the BEV image. The markers between the video frames and those on the BEV image are aligned. Inputting these labeled frames and BEV images into the VLMs led to further improvements in GPT-4o. Unfortunately, these two steps do not lead to improvements for Qwen2-VL. We believe establishing the connection between video frames and global information is useful. However, due to smaller-scale VLMs’ limited visual understanding capabilities, they cannot directly improve their 3D scene comprehension through zero-shot prompts alone. We may need to build a specific dataset and further finetune Qwen2-VL, which could lead to improved performance.
4 Methodology
In this section, we introduce our method. We present the GPT4Scene framework in Section 4.1, which enhances VLMs’ 3D scene understanding using video inputs. In Section 4.2, we discuss how we leverage zero-shot prompts to unlock the potential of powerful closed-source VLMs. In Section 4.3, we apply fine-tuning to enhance smaller-sized open-source VLMs for 3D understanding.
4.1 GPT4Scene Framework
Figure 3 shows GPT4Scene framework and explain how providing global scene information and object-level annotations improves the VLM’s understanding of 3D scenes. First, we assume the captured video is taken while moving around the indoor scene. It consists of $N$ frames, represented as $\bm{V}=\{I_{1},I_{2},\ldots,I_{N}\}$. Processing image sequences with VLMs poses challenges like limited image capacity, rapid context consumption, and higher inference costs. So we uniformly sample $n$ frames:
$s_{i}=[\ (i-1)\cdot N/n\ ]+1,\quad i=1,2,\ldots,n$ (1)
Here, $\{I_{s_{1}},I_{s_{2}},\ldots,I_{s_{n}}\}$ represents the sampled frames with indices $s_{i}$. We denote the sampled video as:
$\bm{V}_{\bm{s}}=\{I_{s_{1}},I_{s_{2}},\ldots,I_{s_{n}}\}$ (2)
This pre-selection dramatically reduces the time and cost of VLMs during training and inference without losing significant indoor scene information.
Global Information: 3D BEV Map. Egocentric video captures only local information, missing the broader scene context. To address this, we reconstruct the entire scene as a point cloud and render a panoramic image as a bird’s-eye view, giving the VLM a complete scene overview. Specifically, starting with an indoor scene video $\bm{V}=\{I_{1},I_{2},\ldots,I_{N}\}$ and the corresponding camera extrinsic $\bm{E}=\{E_{1},E_{2},\ldots,E_{N}\}$, we use 3D reconstruction techniques to generate 3D meshes and point clouds:
$\bm{P}=\mathcal{R}\left(\left\{(I_{n},E_{n})\mid n=1,2,\ldots,N\right\}\right)$ (3)
Here, $\mathcal{R}(\cdot)$ denotes the reconstruction process, and we assume that the camera intrinsics are known. Then, we generate a BEV image of the scene from the global point cloud:
$\bm{I}_{\bm{b}}=\mathcal{T}(\bm{P},E_{t})$ (4)
Here, $E^{t}$ denotes the camera extrinsic for the top-down view, and $\mathcal{T}(\cdot)$ represents the process of rendering the corresponding view based on the camera extrinsic, resulting in a BEV image of the scene. Notably, we continue to provide global 3D information to VLMs in the form of images.
Spatial-Temporal Object Markers. To help VLMs focus on specific objects, we introduce Spatial-Temporal Object markers (STO markers), ensuring consistency between 2D frames and the 3D BEV image. Obtain the 3D point cloud $\bm{P}$ reconstructed from the input video $\bm{V}$. Applying 3D instance segmentation like Mask3D [82] yields instance masks $\{M_{1},M_{2},\ldots M_{K}\}$, where $K$ denotes the total number of objects in the scene.
For BEV images, we first project the 3D mask onto the $xy$-plane and then extract the center coordinates of the bounding box formed by the projection, represented as $(C^{xy}_{1},C^{xy}_{2},\ldots,C^{xy}_{K})$, which are then displayed on the BEV image. For egocentric 2D markers, we first project $\{M_{1},M_{2},\ldots,M_{K}\}$ onto the video frames and then use the center of the bounding box formed by the 2D masks as the 2D markers, represented as $(C^{uv}_{1},C^{uv}_{2},\ldots,C^{uv}_{K})$. The 2D frames with markers and the BEV image are denoted as:
$\bm{V}_{\bm{s}}^{\prime}$ $=\bm{V}_{\bm{s}}(C^{uv}_{1},C^{uv}_{2},\ldots,C^{uv}_{K}),$ (5)
$\bm{I}_{\bm{b}}^{\prime}$ $=\bm{I}_{\bm{b}}(C^{xy}_{1},C^{xy}_{2},\ldots,C^{xy}_{K})$ (6)
Here, $\bm{V}_{\bm{s}}^{\prime}$ and $\bm{I}_{\bm{b}}^{\prime}$ refer to the videos or images with STO markers. We can observe that the 2D and 3D markers are spatially aligned, representing the same object. Additionally, $(C^{uv}_{1},C^{uv}_{2},\ldots,C^{uv}_{K})$ are also consistent across different frames, achieving temporal alignment.
4.2 Unlocking VLMs with Zero-shot Prompts
We evaluate VLMs in a zero-shot setting, initially focusing on powerful closed-source VLMs (e.g., GPT-4o) to assess whether the GPT4Scene framework can effectively enable 3D scene comprehension. This process is called ”unlocking,” which enables VLMs to understand 3D scenes through prompts without additional training. Specifically, we input $\bm{V}_{\bm{s}}^{\prime}$ and $\bm{I}_{\bm{b}}^{\prime}$. To reduce the cost, we stitch the images from $\bm{V}_{\bm{s}}^{\prime}$ into a single large image. We evaluate three tasks: 3D question answering, dense captioning, and visual grounding. In 3D question answering, the goal is to answer scene-related questions like, ”What is the color of the floor?” In dense captioning, the task is to describe a specific object, such as ”Describe the object represented by $C_{5}$.” In visual grounding, the aim is to identify the object ID from a description, like ”What is the ID of the black chair next to the window?” While question answering is independent of object labels, dense captioning, and visual grounding require object markers. These tasks involve detecting objects and filtering based on the IoU of their bounding boxes. Consistent with Chat-Scene [40] and Robin3D [50], we use Mask3D segmentation results as predicted bounding boxes to calculate the IoU.
In addition to traditional tasks, we have conducted further experiments in this zero-shot setting. Qualitative results are shown in Figure 4. By inputting $\bm{V}_{\bm{s}}^{\prime}$ and $\bm{I}_{\bm{b}}^{\prime}$, VLMs can understand the global features of indoor scenes. At this point, GPT-4o can still accept additional first-person perspective frames, allowing it to understand the current position in the scene to plan the following action. Additionally, using GPT-4o as an agent, VLMs can determine the task type based on the given question and choose the appropriate prompt. Therefore, the GPT4Scene framework shows excellent potential as a core technology for the next generation of embodied intelligence.
3D Question Answering (Scene Level) 3D Dense Caption (Object Level) 3D Visual Grounding (Object Level)
ScanQA SQA3D Scan2cap Multi3DRef ScanRefer
26,138 26,623 35,056 41,408 35,061
Total Samples: 164,286
Table 1: Text Annotations of ScanAlign. We obtained the text annotations of ScanAlign by diversifying the text annotations related to ScanNet, resulting in 165K text annotations.
Figure 4: Qualitive Results. We demonstrated GPT-4o’s performance with the zero-shot GPT4Scene prompt, enabling it to understand 3D scenes from video frames and BEV images. It can also accept other images and perform various embodied tasks.
4.3 Enhancing VLMs with ScanAlign Fine-Tuning
Zero-shot prompts can unlock the 3D understanding capabilities of powerful VLMs, but as shown in  Figure 2, this approach does not improve smaller VLMs. Therefore, we aim to enhance open-source, smaller VLMs through fine-tuning. We first construct an indoor scene dataset, ScanAlign, with egocentric, BEV images and text annotations based on the ScanNet [25]. The dataset includes three 3D vision-related tasks represented as $(\bm{V}_{\bm{s}}^{\prime},\bm{I}_{\bm{b}}^{\prime},T)$. The visual input consists of selected video frames with STO markers and BEV images, and $T$ denotes text annotations derived from five ScanNet annotations [25], as shown in Table 1. We use prompts to randomly vary the annotation format to increase annotation diversity, with further details in the supplementary materials. The dataset contains around 165K annotations in total.
Since our method does not require additional modality alignment steps, we can directly perform single-stage instruction fine-tuning on the ScanAlign dataset to enhance the model’s 3D spatial understanding capabilities. During the training phase, the training loss is the Cross-Entropy loss of the language model. The goal is to optimize the learnable parameters, denoted as $\theta$, by minimizing the negative log-likelihood of the target answer, $t^{a}$; we unify the system messages and the user’s question as $t^{q}$. Therefore, the loss function can be expressed as follows:
$\mathcal{L}(\theta)=-\sum_{i=1}^{k}\log P(t^{a}_{i}|t^{a}_{[1,\ldots,i-1]},t^{q}),$ (7)
$k$ represents the number of tokens in the response sequence, and $t^{a}[1,..,i-1]$ denotes the previous $i-1$ tokens in the response. The set of learnable parameters $\theta$ is the vision language projection layers.
After fine-tuning with ScanAlign, during inference, we can input $(\bm{V}_{\bm{s}}^{\prime},\bm{I}_{\bm{b}}^{\prime},Q)$, where $Q$ represents the question. Alternatively, we can inference without the 3D BEV image, using only $(\bm{V}_{\bm{s}}^{\prime},Q)$. For the 3D question answering task, we can even remove all object markers and perform inference using only the original video frames, i.e., $(\bm{V}_{\bm{s}},Q)$, without any additional processing. For the 3D question answering task, we can even remove all object markers, inference with $(\bm{V}_{\bm{s}},Q)$ and use only the original video as input. Our experiments demonstrate that following ScanAlign fine-tuning, small-scale VLMs’ 3D scene understanding capabilities are significantly enhanced.
3D Question Answering Point Encoder Vision Encoder ScanQA (val) SQA3D (val)
Methods BLEU-1 BLEU-4 METEOR ROUGE CIDEr EM-1 EM-R1
Task-Specific Model
ScanQA [6] 30.2 10.1 13.1 33.3 64.9 - -
SQA3D [66] - - - - - 46.6 -
3D-VLP [49] 30.5 11.2 13.5 34.5 - - -
3D-Vista [120] - - 13.9 35.7 - 48.5 -
3D LLM Based Model
Chat-3D [94] 29.1 6.4 11.9 28.5 53.2 - -
Chat-3D v2 [39] 38.4 7.3 16.1 40.1 77.1 - -
3D-LLM [38] 39.3 12.0 14.5 35.7 69.4 - -
LL3DA [21] - 13.5 15.9 37.3 76.8 - -
PQ3D [121] - - - - - 47.1 -
LEO [41] - 11.5 16.2 39.3 80.0 50.0 52.4
Chat-Scene [40] 43.2 14.3 18.0 41.6 87.7 54.6 57.5
Vision LLM Based Model  (Close-Source)
GPT-4o [71] 30.9 10.6 15.2 32.6 66.0 40.3 46.1
GPT-4o (GPT4Scene) 35.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+4.7}}}$ $\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+2.1}}}$ 15.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.4}}}$ 37.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+5.1}}}$ 73.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+7.5}}}$ 42.8$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+2.5}}}$ 49.4+3.$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+3.3}}}$
Vision LLM Based Model  (Open-Source)
InternVL2-8B [23] 23.9 3.3 14.5 34.3 62.5 33.0 45.3
MiniCPM-V-2.6 [105] 25.1 8.4 11.8 31.5 60.1 42.6 46.6
Qwen2-VL-7B [90] 27.8 3.0 11.4 29.3 53.9 40.7 46.7
Qwen2-VL-7B (GPT4Scene) 43.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+15.6}}}$ 14.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+11.6}}}$ 17.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+6.3}}}$ 43.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.2}}}$ 90.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+37.0}}}$ 57.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.7}}}$ 60.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.1}}}$
Qwen2-VL-7B (GPT4Scene)-HD 41.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.1}}}$ 15.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+12.9}}}$ 17.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+6.2}}}$ 43.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.2}}}$ 89.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+36.0}}}$ 57.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.5}}}$ 60.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+13.5}}}$
Qwen2-VL-7B (GPT4Scene)-HDM 44.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.6}}}$ 15.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+12.5}}}$ 18.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+7.5}}}$ 46.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+17.1}}}$ 96.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+42.4}}}$ 59.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+18.7}}}$ 62.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+15.5}}}$
Table 2: Evaluation of 3D Question Answering on ScanQA [6] and SQA3D [66] datasets. GPT-4o (GPT4Scene) in the zero-shot setting outperforms most 3D LLM models. Fine-tuned with GPT4Scene, Qwen2-VL achieves state-of-the-art performance. The base setting uses $N=8$ frames at $128\times 123$, ”HD” increases resolution to $512\times 490$, and ”HDM” combines this resolution with $N=32$ frames.
5 Experiments
In this section, we present the experimental results.  Section 5.1 outlines the implementation details, while  Section 5.2 provides the main results for 3D question answering, dense captioning, visual grounding, and qualitative results. Finally, Section 5.3 covers the ablation study, demonstrating that including global information and STO markers is effective in the training process.
5.1 Implantation Details
All our scene data is based on the ScanNet dataset [25], which contains 1,513 scenes. All these benchmark datasets use the same data split as ScanNet. Our benchmark, as introduced in  Table 1, is divided into three tasks. For the 3D question answering task, we adopt ScanQA [6] and SQA3D [66] as benchmarks. For the 3D dense captioning task, we select Scan2Cap as the benchmark [22]. For the 3D visual grounding task, we use ScanRefer [13] and Multi3DRef [114] to evaluate grounding capabilities. Compared to ScanRefer [13], Multi3DRef evaluates multiple objects, making it more comprehensive and reasonable. We use the same metrics as the original version.
We use the original reconstruction method [26] from ScanNet for reconstruction and align the point cloud. We then use Mask3D [82] to do the 3D instance segmentation, which is consistent with Chat-Scene [40] and Robin3D [50]. After obtaining the 3D segmentation result, we take the xy-coordinates of its center point as the marker’s location in the BEV image. Next, we project the mask onto 2D frames and use the center point of the bounding box of the 2D mask as the marker position. We can directly input the original video during inference to process question answering tasks. However, the markers on 2D images remain essential for dense captioning and visual grounding tasks due to the task’s requirements.
We select $N=8$ frames for each video with a $128\times 123$ resolution. For closed-source VLMs, we use GPT-4o [71], which excels at single images over multiple frames. So, we stitch 8 frames into a large image. For open-source VLMs, the sampled frames are input directly. We use Qwen2-VL-7B-Instruct [90, 117], and in addition to the base setting, we introduce an HD setting (”high resolution”), where the resolution is $512\times 490$. Furthermore, we introduce an HDM setting (”high resolution, more frames”), where the resolution is also $512\times 490$, but with $N=32$ frames. All parameters are fine-tuned using a base learning rate of $5\times 10^{-6}$, cosine annealing, a batch size of 128, and one epoch (1283 steps, 0.1 warmup ratio). Training took approximately six hours on 8 NVIDIA A100 GPUs.
5.2 Main Results
3D Question Answering. The 3D question answering results are shown in Table 2. We categorize the methods into three types: classic task-specific models focused on 3D QA tasks, 3D point LLM-based models, and vision LLM-based models. GPT-4o (GPT4Scene), in a zero-shot manner, outperforms all task-specific models, highlighting the effectiveness of GPT4Scene as the prompt. In contrast, open-source VLMs perform poorly in zero-shot mode without fine-tuning with GPT4Scene, consistent with our pilot study findings. Using the strategy proposed by GPT4Scene, we fine-tune Qwen2-VL-7B, achieving superior question answering results. Notably, Qwen2-VL-7B (GPT4Scene) outperforms all other methods, reaching state-of-the-art performance. Furthermore, compared to the original Qwen2-VL-7B, BLEU-1 in ScanQA improved by 56.1% (27.8 $\xrightarrow{}$ 43.4), and CIDEr increased by 68.6% (53.9 $\xrightarrow{}$ 90.9). In SQA3D, EM-1 rose by 41.0% (40.7 $\xrightarrow{}$ 57.4). Relative to GPT-4o in zero-shot mode, these metrics show gains of 21.9%, 23.7%, and 34.1%, respectively. Our approach greatly improves the model’s understanding of 3D indoor scenes.
3D Dense Caption & Visual Grounding. We also evaluate the model’s performance in dense captioning and visual grounding, which differ from question answering as they require markers to accomplish these tasks. The experimental results are presented in Table 3 and Table 4. As shown, Qwen2-VL-7B, fine-tuned with GPT4Scene, significantly improves VLMs’ 3D captioning and grounding capabilities. Additionally, under high-resolution and larger frame settings (HD), our model achieves state-of-the-art (SOTA) performance, surpassing all existing methods.
Qualitative Results. We present qualitative results in Figure 4, conducted in a zero-shot setting on GPT-4o. Stitched frames provide an overview of the scene, while individual frames capture details and actions. Beyond standard tasks like object captioning, spatial description, and counting, GPT4Scene handles embodied tasks, such as directing users to retrieve paper from a nearby bookshelf. In the last row, pink-highlighted segments indicate BEV images as input, enhancing navigation capabilities. GPT4Scene also excels in navigation and patrolling tasks, performing industrial inspections by observing machine indicators.
3D Dense Caption IoU@0.25 IoU@0.5
Methods BLEU-4 ROUGE BLEU-4 ROUGE
Task-Specific Model
Scan2Cap [22] 34.2 55.3 23.3 44.5
3DJCG [9] 40.2 59.2 31.0 50.8
X-Trans2Cap [108] 35.7 54.7 25.1 45.3
3D-VisTA [120] 36.5 57.6 34.0 54.3
Vote2Cap-DETR [20] 39.3 59.3 34.5 54.4
3D LLM Based Model
LL3DA [21] 41.4 59.5 36.8 55.1
LEO [41] 36.9 57.8
Chat-Scene [40] 38.2 60.6 36.3 58.1
Robin3D [50] 38.4
Vision LLM Based Model
Qwen2-VL-7B [90] 3.8 24.7 3.8 24.6
Qwen2-VL-7B (GPT4Scene) 36.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.5}}}$ 57.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.9}}}$ 34.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+30.4}}}$ 55.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+30.6}}}$
Qwen2-VL-7B (GPT4Scene)-HD 40.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+36.6}}}$ 60.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+35.5}}}$ 37.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+34.1}}}$ 57.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+33.1}}}$
Qwen2-VL-7B (GPT4Scene)-HDM 43.1$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+39.3}}}$ 61.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+37.2}}}$ 40.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+36.8}}}$ 59.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+34.7}}}$
Table 3: Evaluation of 3D Dense Caption on Scan2Cap [22]. Our results outperform those of existing 3D LLM based models.
5.3 Ablation Study
In this section, we conduct ablation studies to validate our design modules and analyze factors like resolution and the number of frames.
Ablation on Our Modules. Figure 2 demonstrates that BEV images and STO markers enhance spatial understanding. We use the 3D question answering (QA) task to validate this further because pure video input without STO markers or BEV images can suffice. At the same time, markers are essential for 3D dense caption and visual grounding to reference objects during evaluation. As shown in Table 5, removing BEV images during training and inference reduces performance on both tasks. Further removal of STO markers causes an additional drop in QA performance, highlighting the critical role of BEV images and STO markers in enabling VLMs to understand 3D scenes.
Ablation on Frames and Resolution. Here, we conduct ablation experiments on additional factors using 3D question answering and visual grounding as benchmarks. The results are shown in Table 6. The first three rows show that image resolution significantly affects visual grounding performance but has limited improvement on the QA task. Lastly, the results in the final three rows indicate that increasing the number of frames enhances indoor scene understanding, with a more pronounced effect on grounding performance compared to the limited improvement in QA.
3D Visual Grounding ScanRefer Multi3DRef
Methods Acc@0.25 Acc@0.50 all F1@0.25 all F1@0.50
Task-Specific Model
3DVG-Transformer [115] 47.6 34.7 25.5
3DJCG [9] 49.6 37.3 26.6
D3Net [14] 37.9 32.2
M3DRef-CLIP [114] 51.9 44.7 42.8 38.4
3D LLM Based Model
Chat-Scene [40] 55.5 50.2 57.1 52.4
Vision LLM Based Model
Qwen2-VL-7B [90] 5.4 5.1 21.1 19.9
Qwen2-VL-7B (GPT4Scene) 40.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+35.1}}}$ 36.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+31.6}}}$ 45.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+24.3}}}$ 42.1+22.$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+22.2}}}$
Qwen2-VL-7B (GPT4Scene)-HD 50.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+45.5}}}$ 46.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+41.3}}}$ 53.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.6}}}$ 50.0$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+30.1}}}$
Qwen2-VL-7B (GPT4Scene)-HDM 62.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+51.9}}}$ 57.0$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.3}}}$ 64.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+43.4}}}$ 59.8$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+39.9}}}$
Table 4: Evaluation of 3D Visual Grounding on ScanRefer [13] and Multi3DRef [114]. Our method reaches SOTA performance.
Ablation ScanQA Multi3DRef@0.5
ROUGE CIDEr MT ALL
Qwen2-VL-7B (GPT4Scene) 43.6 90.9 36.3 42.1
w/o BEV Image 42.3 87.1 27.8 32.1
w/o BEV Image & STO markers 41.7 85.0 - -
Table 5: Ablation Study. Removing BEV images during training and inference leads to a performance decrease, and further removal of STO markers causes an additional performance decrease.
Num Resolution ScanQA ScanRefer
Frames ROUGE CIDEr Acc0.25 Acc0.5
8 128 (base) 43.6 90.9 40.5 36.7
256 43.8 90.0 49.2 44.8
512 (HD) 43.6 89.9 50.9 46.4
16 512 45.4 93.4 58.6 53.4
32 512 (HDM) 46.5 96.3 62.6 57.0
Table 6: Ablation Study on frames and resolution. Results show that the number of frames affects both QA and grounding, while resolution significantly impacts grounding.
6 Conclusion
We introduce GPT4Scene, a framework that enhances Vision-Language Models (VLMs) to comprehend 3D scenes directly from pure vision input. Our pilot study reveals that global scene information and the correspondence between video frames and objects in the global context are essential for improving the 3D understanding of VLMs. We propose reconstructing a 3D point cloud from input video to generate a BEV (Bird’s Eye View) image for global information. We establish global and local data correspondences by adding spatial-temporal object markers (STO markers) to both video frames and BEV images. GPT4Scene achieves excellent performance with closed-source VLMs like GPT-4o using zero-shot inference. For smaller VLMs, such as Qwen2-VL, we create the ScanAlign dataset, which includes video frames, BEV images, STO markers, and 165K text annotations. After fine-tuning, VLMs show significant improvements in scene understanding, reaching state-of-the-art performance in question-answering tasks. Moreover, fine-tuned VLMs can perform well on QA tasks with only original video frames, demonstrating that GPT4Scene enables VLMs to understand 3D scenes effectively.
7 Appendix
Appendix A Prompts of Closed-source VLMs
Here, we present the prompts used by GPT4Scene in the closed-source VLMs (GPT-4o), as illustrated in Figure 5. The process begins with a general system prompt, which outlines the overview of two images provided as input. The first image is a stitched 2D view captured from a video, with dimensions of $2\times 4$. The second image represents a BEV (Bird’s Eye View). Subsequently, we perform evaluations across various tasks and benchmarks, with each benchmark associated with a specific prompt. We take the 3D question-answering task on ScanQA as an example. The benchmark prompt consists of three parts:
  1. 1.
    Important Guidelines: It clarifies that while we provide object IDs for reference, they cannot be directly used when answering questions. Additionally, it specifies adapting the response style to match that of ScanQA. Since ScanQA’s responses are typically short single words, we aim to keep the answers concise within the benchmark prompt, targeting 1-5 words.
  2. 2.
    Answer Format: In this part, we use a standardized regularized format to structure the answers, which helps improve accuracy when addressing questions.
  3. 3.
    Examples: we include two example cases.
Our zero-shot prompting process is illustrated in the bottom-left corner Figure 5. The system prompt and ScanQA prompt are used as the system message. In the user message, we input the stitched image and BEV image. Finally, the query message includes the question. The responses generated through this process require further refinement, as depicted in the bottom-right corner of Figure 5. First, we remove the regularized formatting from the answers. Next, we clean the answers by addressing singular/plural forms and case sensitivity. This final step ensures that we obtain the refined answers.
Task type Model
Qwen2-VL Ours
Action Sequence 85.5 82.0
Action Prediction 69.5 70.5
Action Antonym 83.0 86.0
Fine-grained Action 51.5 51.5
Unexpected Action 82.0 78.0
Object Existence 87.5 88.5
Object Interaction 82.0 81.5
Object Shuffle 41.0 45.0
Moving Direction 42.0 40.0
Action Localization 65.0 66.5
Scene Transition 93.5 94.0
Action Count 47.5 43.5
Moving Count 69.5 71.5
Moving Attribute 90.0 88.5
State Change 48.0 49.0
Fine-grained Pose 63.0 63.5
Character Order 74.5 71.0
Egocentric Navigation 39.5 41.5
Episodic Reasoning 47.0 47.0
Counterfactual Inference 62.5 65.5
Avg 66.2 66.225
Table 7: The result of MVBench [57]. After fine-tuning with ScanAlign in GPT4Scene, our model shows improved 2D understanding, particularly in object and action metrics.
      Benchmark       Model
      Qwen2-VL       Ours
      MMBench-ENval [109]       82.4       81.2
      MMBench-CNval [109]       81.7       79.9
      MMStar [17]       60.7       57.6
      RealWorldQA [98]       70.1       68.5
      Video-MME [31]       59.8       58.4
Table 8: The result of 2D Multi-modal Benchmark. After fine-tuning with ScanAlign in GPT4Scene, our model’s 2D understanding capabilities did not decline.
Methods BLEU-n Metrics Language Generation Metrics
EM-1 BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE METEOR CIDEr
Task-Specific Model
ScanQA [6] 21.1 30.2 20.4 15.1 10.1 33.3 13.1 64.9
3D-VLP [49] 21.7 30.5 21.3 16.7 11.2 34.5 13.5 67.0
3D-Vista [120] 13.9 35.7
3D LLM Based Model
3D-LLM [38] 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
LL3DA [21] 13.5 37.3 15.9 76.8
LEO [41] 24.5 11.5 39.3 16.2 80.0
Scene-LLM [32] 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
Chat-Scene [40] 21.6 43.2 29.1 20.6 14.3 41.6 18.0 87.7
Vision LLM Based Model
InternVL2-8B [23] 19.0 23.9 12.0 6.3 3.3 34.3 14.5 62.5
MiniCPM-V-2.6 [105] 20.6 25.1 16.9 12.4 8.4 31.5 11.8 60.1
Qwen2-VL-7B [90] 19.0 27.8 13.6 6.3 3.0 29.3 11.4 53.9
Qwen2-VL-7B (GPT4Scene) 25.5 43.4 29.3 20.9 14.6 43.6 17.7 90.9
Qwen2-VL-7B (GPT4Scene)-HD 26.3 41.9 28.6 21.6 15.9 43.6 17.6 89.9
Qwen2-VL-7B (GPT4Scene)-HDM 28.2 44.4 30.3 22.3 15.5 46.5 18.9 96.3
Table 9: Full Evaluation of 3D Question Answering on ScanQA [6].
      Methods       Test Set       Avg.(EM-1)       EM-R1
      What       Is       How       Can       Which       Others
      Task-Specific Model
      ClipBERT [66]       30.2       60.1       38.7       63.3       42.5       42.7       43.3       –
      SQA3D [66]       31.6       63.8       46.0       69.5       43.9       45.3       46.6       –
      3D-VisTA [120]       34.8       63.3       45.4       69.8       47.2       48.1       48.5       –
      3D LLM Based Model
      PQ3D [121]       37.1       61.3       44.5       60.9       47.0       45.1       47.1       –
      LEO [41]       –       –       –       –       –       –       50.0       52.4
      Scene-LLM [32]       40.9       69.1       45.0       70.8       47.2       52.3       54.2       –
      Chat-Scene [40]       45.4       67.0       52.0       69.5       49.9       55.0       54.6       57.5
      Vision LLM Based Model
      InternVL2-8B [23]       30.5       53.8       5.5       47.3       25.8       36.3       33.0       45.3
      MiniCPM-V-2.6 [105]       34.8       55.2       33.0       49.8       43.3       48.3       42.6       46.6
      Qwen2-VL-7B [90]       29.0       59.2       33.4       50.5       44.2       43.2       40.7       46.7
      Qwen2-VL-7B (GPT4Scene)       50.7       70.9       48.0       70.5       52.9       59.3       57.4       60.7
      Qwen2-VL-7B (GPT4Scene)-HD       51.4       69.1       50.2       69.4       51.3       57.9       57.2       60.4
      Qwen2-VL-7B (GPT4Scene)-HDM       55.9       69.9       50.8       68.7       53.3       60.4       59.4       62.4
Table 10: Full Evaluation of 3D Question Answering on SQA3D [66].
Appendix B 2D Multi-modal Benchmark
We tested the fine-tuned Qwen2-VL-7B(GPT4Scene) model on 2D image and video multimodal large models. Table 7 shows the results of MVBench. As we can see, our model shows improvement for the object and action metrics, indicating that the model fine-tuned with ScanAlign is better at handling spatial variations and the information of objects in the scene. Table 8 presents the results on other benchmarks, where we can observe that after our fine-tuning, the model’s ability to understand images and videos did not decline significantly. It demonstrates the effectiveness of ScanAlign.
Appendix C Qualitative Results
Figures 6, 7, 8 and 9 presents our qualitative results, derived from Qwen2-VL after fine-tuning on ScanAlign. Figures 6 and 7 showcase the results of 3D question answering, where only unmarked videos were provided as input. We observe that the model can answer questions from ScanQA [6] and SQA3D [66]. Meanwhile, Figures 8 and 9 illustrate the model’s performance on 3D dense captioning and 3D visual grounding tasks, where marked inputs were required. These results demonstrate that our model can provide accurate responses, relying solely on visual information without requiring 3D point cloud input.
Appendix D Full Quantitive Results
Here, we present the complete metrics for all five benchmarks. Table 9 and Table 10 show results for ScanQA [6] and SQA3D [66]. Table 11 provides the full results for Scan2Cap [22], while Table 12 and Table 13 present the results for ScanRefer [13] and Multi3DRef [114]. Our model significantly improves across all benchmarks, highlighting that only pure vision language models can understand 3D scenes effectively.
Methods IoU@0.25 IoU@0.5
CIDEr BLEU-4 METEOR ROUGE CIDEr BLEU-4 METEOR ROUGE
Task-Specific Model
Scan2Cap [22] 56.8 34.2 26.3 55.3 39.1 23.3 22.0 44.5
3DJCG [9] 64.7 40.2 27.7 59.2 49.5 31.0 24.2 50.8
X-Trans2Cap [108] 61.8 35.7 26.6 54.7 43.9 25.1 22.5 45.3
D3Net [14] 62.6 35.7 25.7 53.9
3D-VLP [49] 70.7 41.0 28.1 59.7 54.9 32.3 24.8 51.5
Vote2Cap-DETR [20] 71.5 39.3 28.3 59.3 62.6 35.7 25.7 53.9
3D-VisTA [120] 71.0 36.5 28.4 57.6 66.9 34.0 27.1 54.3
3D LLM Based Model
LL3DA [21] 74.2 41.4 27.8 59.5 65.2 36.8 26.0 55.1
LEO [41] 68.4 36.9 27.7 57.8
Chat-Scene [40] 81.9 38.2 29.0 60.6 77.2 36.3 28.0 58.1
Robin3D [50] 87.2 38.4
Vision LLM Based Model
Qwen2-VL-7B [90] 0.0 3.8 16.7 24.7 0.0 3.8 16.5 24.6
Qwen2-VL-7B (GPT4Scene) 63.8 36.3 26.5 57.6 60.6 34.2 25.6 55.2
Qwen2-VL-7B (GPT4Scene)-HD 79.1 40.4 28.3 60.2 74.4 37.9 27.3 57.7
Qwen2-VL-7B (GPT4Scene)-HDM 91.7 43.1 29.3 61.9 86.3 40.6 28.2 59.3
Table 11: Full Evaluation of 3D Dense Caption on Scan2Cap [22].
Methods Unique Multiple Overall
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
Task-Specific Model
ScanRefer [13] 76.3 53.5 32.7 21.1 41.2 27.4
TGNN [42] 68.6 56.8 29.8 23.2 37.4 29.7
SAT [108] 73.2 50.8 37.6 25.2 44.5 30.1
InstanceRefer [107] 75.7 64.7 29.4 23.0 38.4 31.1
3DVG-Transformer [115] 81.9 60.6 39.3 28.4 47.6 34.7
MVT [43] 77.7 66.4 31.9 25.3 40.8 33.3
3D-SPS [64] 84.1 66.7 40.3 29.8 48.8 37.0
ViL3DRel [19] 81.6 68.6 40.3 30.7 47.9 37.7
3DJCG [9] 83.5 64.3 41.4 30.8 49.6 37.3
D3Net [14] 72.0 30.1 37.9
BUTD-DETR [46] 84.2 66.3 46.6 35.1 52.2 39.8
HAM [15] 79.2 67.9 41.5 34.0 48.8 40.6
3DRP-Net [93] 83.1 67.7 42.1 32.0 50.1 38.9
3D-VLP [49] 84.2 64.6 43.5 33.4 51.4 39.5
EDA [97] 85.8 68.6 49.1 37.6 54.6 42.3
M3DRef-CLIP [114] 85.3 77.2 43.8 36.8 51.9 44.7
3D-VisTA [120] 81.6 75.1 43.7 39.1 50.6 45.8
ConcreteNet [87] 86.4 82.1 42.4 38.4 50.6 46.5
DOrA [95] 52.8 44.8
3D LLM Based Model
Chat-Scene [40] 89.6 82.5 47.8 42.9 55.5 50.2
Robin3D [50] 60.8 55.1
Vision LLM Based Model
Qwen2-VL-7B [90] 6.3 6.3 5.1 4.8 5.4 5.1
Qwen2-VL-7B (GPT4Scene) 65.5 61.2 34.8 31.1 40.5 36.7
Qwen2-VL-7B (GPT4Scene)-HD 77.5 71.9 44.9 40.6 50.9 46.4
Qwen2-VL-7B (GPT4Scene)-HDM 90.3 83.7 56.4 50.9 62.6 57.0
Table 12: Full Evaluation of 3D Visual Grounding on ScanRefer [13].
Methods ZT w/o D ZT w/ D ST w/o D ST w/ D MT ALL
F1 F1 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5
Task-Specific Model
3DVG-Trans+ [115] 87.1 45.8 27.5 16.7 26.5 25.5
D3Net (Grounding) [14] 81.6 32.5 38.6 23.3 35.0 32.2
3DJCG (Grounding) [9] 94.1 66.9 26.0 16.7 26.2 26.6
M3DRef-CLIP [114] 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
3D LLM Based Model
Chat-Scene [40] 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
Vision LLM Based Model
Qwen2-VL-7B [90] 84.8 58.5 20.1 19.1 14.7 13.5 16.8 15.4 21.1 19.9
Qwen2-VL-7B (GPT4Scene) 85.2 61.4 60.1 55.1 37.7 34.4 39.4 36.3 45.4 42.1
Qwen2-VL-7B (GPT4Scene)-HD 93.6 81.8 72.5 66.2 46.6 42.9 41.8 38.9 53.7 50.0
Qwen2-VL-7B (GPT4Scene)-HDM 97.4 84.4 85.0 77.7 59.9 55.1 48.6 44.6 64.5 59.8
Table 13: Full Evaluation of 3D Visual Grounding on Multi3DRef [114].
Figure 5: Prompts of Closed-sourse VLMs. We show the prompts used for GPT-4o (GPT4Scene), which consist of a system prompt and a benchmark prompt. After generating responses, we further refine them.
Figure 6: Qualitive Results: Question Answering. We provide videos without object annotations.
Figure 7: Qualitive Results: Question Answering. We provide videos without object annotations.
Figure 8: Qualitive Results: 3D Dense Caption and Visual Grounding. We provide videos with object annotations.
Figure 9: Qualitive Results: 3D Dense Caption and Visual Grounding. We provide videos with object annotations.