### v4

# GPT4Scene: Understand 3D Scenes from Videos with Vision-Language Models

Abstract
In recent years, 2D Vision-Language Models (VLMs) have made significant strides in image-text understanding tasks. However, their performance in 3D spatial comprehension, which is critical for embodied intelligence, remains limited. Recent advances have leveraged 3D point clouds and multi-view images as inputs, yielding promising results. However, we propose exploring a purely vision-based solution inspired by human perception, which merely relies on visual cues for 3D spatial understanding. This paper empirically investigates the limitations of VLMs in 3D spatial knowledge, revealing that their primary shortcoming lies in the lack of global-local correspondence between the scene and individual frames. To address this, we introduce GPT4Scene, a novel visual prompting paradigm in VLM training and inference that helps build the global-local relationship, significantly improving the 3D spatial understanding of indoor scenes. Specifically, GPT4Scene constructs a Bird’s Eye View (BEV) image from the video and marks consistent object IDs across both frames and the BEV image. The model then inputs the concatenated BEV image and video frames with markers. In zero-shot evaluations, GPT4Scene improves performance over closed-source VLMs like GPT-4o. Additionally, we prepare a processed video dataset consisting of 165K text annotation to fine-tune open-source VLMs, achieving state-of-the-art performance on all 3D understanding tasks. Surprisingly, after training with the GPT4Scene paradigm, VLMs consistently improve during inference, even without object marker prompting and BEV image as explicit correspondence. It demonstrates that the proposed paradigm helps VLMs develop an intrinsic ability to understand 3D scenes, which paves the way for a seamless approach to extending pre-trained VLMs for 3D scene understanding.
Figure 1: The overall architecture of GPT4Scene. It is capable of understanding 3D scenes and performing tasks such as 3D question answering, dense captioning, and visual grounding using only video input. In contrast to 3D point LLMs, GPT4Scene takes input solely from the vision modality, with global information provided by the BEV image reconstructed from the 3D structure derived from the video.
1 Introduction
3D scene understanding aims to comprehend the overall layout of the surrounding complex environments and the spatial relationships between objects [5, 21, 12]. It plays a crucial role in applications such as embodied intelligence, virtual reality, and smart cities [40, 20, 9]. With the rapid development of Large Language Models [91, 89, 71, 92, 90], Vision-Language Models have demonstrated impressive performance in image and video understanding [60, 61, 53, 126, 96]. Researchers have extended this paradigm to 3D perception by incorporating point clouds, aiming to improve scene understanding [37, 20, 100, 31, 68, 40, 39, 49].
Recent 3D point LLMs leverage point clouds aligned with LLM features for indoor scene understanding [20, 37]. While combining point clouds and images [128, 20, 39] in Point-Vision-LLM paradigms improves performance through richer visual cues, aligning these modalities with text remains challenging. This complexity motivates exploration of vision-only solutions, inspired by humans’ natural ability to perceive 3D environments without point clouds, offering a promising direction for efficient scene analysis.
In this work, we aim to leverage pre-trained VLMs without modifying their architecture, maximizing their vision perception capabilities. However, their effectiveness in understanding immersive 3D indoor scenes remains limited. Our analysis shows that directly inputting scene videos into VLMs fails in 3D scene understanding due to two factors: i) the lack of a global scene representation, ii) misalignment between per-frame local observations and their spatial-temporal context.
To address this, we propose GPT4Scene, a framework that enhances VLMs’ spatial understanding (see Figure 1). We first perform 3D reconstruction on input videos to generate a Bird’s Eye View (BEV) image, offering a comprehensive scene layout. Additionally, we introduce Spatial-Temporal Object markers (STO-markers) in both the BEV image and 2D frames. These markers maintain consistent object IDs across frames (temporal level) and align with the BEV image (spatial level), bridging the global-local relationship. Empirical results show that GPT4Scene remains robust to reconstruction quality and marker accuracy, as it prioritizes learning global-local correspondences over precise geometric reconstructions.
We first explored the effectiveness of GPT4Scene under a training-free approach. Experimental results revealed that it was notably effective for powerful large-scale VLMs such as Qwen2-VL-72B [96], as well as closed-source models like GPT-4o [72] and Gemini-1.5-Pro [88], significantly enhancing their 3D scene understanding capabilities, even reaching levels comparable to previous state-of-the-art point-based methods. However, the improvements were limited on smaller-scale models. For smaller open-source vision-language models (VLMs), we introduce ScanAlign, a multimodal dataset comprising 165K aligned data pairs featuring STO-marker-annotated video frames, BEV images, and textual descriptions. Fine-tuning Qwen2-VL-7B on ScanAlign achieves state-of-the-art performance, delivering a 48% relative improvement (40.7 → 60.7 EM-1 score) on 3D question answering (SQA3D), surpassing the previous SOTA Chat-Scene by 11.0% (54.6 vs 60.7). The model demonstrates even stronger 3D visual grounding capabilities on Multi3DRef, outperforming Chat-Scene by 13.0% (57.1 → 64.5). These advancements underscore GPT4Scene’s effectiveness in seamlessly enhancing VLMs with strong 3D spatial understanding.
Our paper makes these major contributions:
  • We introduce GPT4Scene, a framework that enhances Vision-Language Models (VLMs) to comprehend 3D scenes directly from pure vision input.
  • We introduce two techniques: i) A 3D BEV image with global context information, and ii) Spatial-Temporal Object markers (STO-markers) for spatial and temporal consistency across BEV image and video frames.
  • We introduce ScanAlign, a dataset comprising video frames, BEV images with STO-markers, and text annotations. Fine-tuning VLMs on this dataset significantly improves 3D understanding, even with raw video inputs.
  • GPT4Scene demonstrates robust performance in zero-shot and fine-tuning settings, achieving SOTA results across diverse 3D scene understanding tasks.
2 Related Work
3D Indoor Scene Understanding. 3D indoor scene understanding allows robots to identify object positions, structures, and relationships within indoor environments, enabling question-and-answer interactions about the scene’s content. This process combines 3D perception with large language models (LLMs). 3D perception, as a foundational component, is typically trained on common indoor datasets [24, 7, 10, 69, 79, 95, 85, 27, 51, 124] using point clouds as input, supporting tasks like 3D object detection and instance segmentation [75, 47, 70, 83, 94, 102, 116]. Recent advancements in 3D Vision-Language Learning (3D-VL) combine 3D perception tasks with natural language, introducing diverse textual annotations on datasets like ScanNet to support tasks such as 3D Question Answering [5, 66, 112, 17], 3D Dense Captioning [21], and 3D Visual Grounding [12, 121, 2, 1, 50]. Initial studies focus on single 3D-VL tasks [114, 19, 103, 6, 35, 43, 45, 64, 110, 122, 32], while recent research introduces unified models for multiple tasks [8, 13]. 2D vision-language pretraining (2D-VLP) has driven progress in 3D visual-language learning (3D-VL) [34, 36, 73, 87, 107, 118, 119, 105], with recent 3D-VLP [127, 128, 48, 28, 108, 46, 97] methods demonstrating that combining 2D visual cues with 3D point clouds enables complementary cross-modal alignment. This multimodal approach helps overcome geometric complexity and sparse annotation challenges in pure point cloud processing.
3D Point Cloud LLMs. 3D vision-language tasks aim to integrate 3D scene understanding with natural language processing. However, we aspire to go further by incorporating 3D content into large language models (LLMs) to achieve more natural human-computer interaction. Initially, this began with 3D point cloud LLMs. 3D point cloud LLMs take point clouds as input, enabling natural language generation and interaction in 3D scenes. Early 3D LLMs focused on object-level geometry and appearance [33, 77, 106, 76]. Later, they expanded to indoor scenes, emphasizing spatial relationships among objects and overall scene features, often utilizing scene point clouds augmented with auxiliary 2D multi-view images [37, 20, 100, 31, 68]. To better capture object relationships, recent 3D LLMs decouple scene objects before feeding them into LLMs [40, 39]. Some approaches rely more heavily on visual inputs to determine scene context [9, 125, 59]. Here, we aim to explore whether pure visual inputs can better handle indoor scene understanding.
Vision Language Models (VLMs). Vision Language Models (VLMs) are multimodal models that integrate visual and language processing capabilities, enabling the understanding and generation of combined image-text information. The origin of VLMs can be traced back to CLIP’s 2D image-text pair pretraining [78, 11, 82], which laid the foundation for incorporating LLMs. Early VLMs used attention mechanisms or Q-Former to fuse image and text modalities before inputting them to LLMs [55, 56, 26, 3, 44]. Later, an approach emerged that directly projects image features into the LLM using an MLP [60, 61, 53, 126, 96], achieving better performance. Building on this, VLMs expanded into visual grounding tasks [98, 74, 80, 52, 29, 86, 15] and further to video understanding by using spatiotemporal compression to process information from long image sequences [117, 23, 67, 65, 4, 58, 84, 81, 62]. Currently, while some studies employ VLMs for indoor scene understanding, most remain either benchmark-oriented [109, 54] or still rely on 3D-based methods [120, 123]. This fundamentally reveals that VLMs cannot directly comprehend 3D scenes, making our core mission to empower them with 3D world interpretation capabilities.
3 Methodology
Figure 2: The Framework of GPT4Scene. A scene video is processed by sampling frames, reconstructing a point cloud, and generating a BEV image. Object locations are detected from the point cloud and projected onto the video frames. The resulting frames and BEV image, enhanced with STO-markers, are inputs for Large Language Model (VLM) training and inference.
In this section, we introduce the GPT4Scene framework as shown in Section 3.1, which enhances VLMs’ 3D scene understanding using video inputs. In Section 3.2, we discuss how we leverage zero-shot prompts to unlock the potential of powerful large-scale VLMs. In Section 3.3, we apply fine-tuning to enhance VLMs for 3D understanding.
3.1 GPT4Scene Framework
This section first presents an overview of the GPT4Scene framework, then details the 3D BEV Map (providing global scene layouts) and STO-markers (delivering localized object cues) as core components.
Overview. Here we introduce the architecture of GPT4Scene. Figure 2 shows how GPT4Scene integrates global scene layouts and object-level details to improve VLMs’ 3D scene understanding. Given a video $\mathcal{V}\!=\!\{I_{1},\ldots,I_{N}\}$ captured by moving around an indoor scene, we first approximately uniformly sample $n$ frames using indices: $s_{i}=\left\lfloor(i-1)\frac{N}{n}\right\rfloor+1,\quad\forall i\in\{1,\ldots,n\}$ to form $\mathcal{V}^{*}\!=\!\{I_{s_{1}},\ldots,I_{s_{n}}\}$. This strategy reduces token counts and computational overhead for Vision-Language Models (VLMs) while preserving scene context, without significant information loss. We then leverage the complete temporal sequence for 3D scene reconstruction, generating a holistic Bird’s-Eye View (BEV) map. Subsequent 3D instance segmentation enables precise object localization, which is projected onto both the BEV map and 2D video frames to establish Spatio-Temporal Object markers (STO-markers). The technical details are elaborated as follows.
Global Information: 3D BEV Map. Egocentric videos lack global scene context. We address this by reconstructing the 3D scene into a point cloud and rendering a bird’s-eye view (BEV) image for holistic VLM understanding. Given video $\mathcal{V}=\{I_{1},I_{2},\ldots,I_{N}\}$ with camera extrinsics =$\mathcal{E}=\{E_{1},E_{2},\ldots,E_{N}\}$, we use 3D reconstruction techniques to generate 3D meshes and point clouds:
$\mathcal{P}=\mathcal{R}\left(\{(I_{t},E_{t})\}_{t=1}^{N}\right)$ (1)
Here, $\mathcal{R}(\cdot)$ denotes the reconstruction process, and we assume that the camera intrinsics are known. Then, we generate a BEV image of the scene from the global point cloud:
$\mathcal{I}_{b}=\mathcal{T}(\mathcal{P},\ E_{top})$ (2)
Here, $E_{\text{top}}\in\text{SE}(3)$ denotes the camera extrinsic for the top-down view, and $\mathcal{T}(\cdot)$ represents the process of rendering the corresponding view based on the camera extrinsic, resulting in a BEV image of the scene. In particular, we continue to provide global 3D information to VLMs in the form of the top-view images instead of points.
Local Correspondence: STO-markers. To help VLMs focus on specific objects, we introduce Spatial-Temporal Object markers (STO-markers), ensuring consistency between 2D frames and the 3D BEV image. Obtain the 3D point cloud $\mathcal{P}$ reconstructed from the input video $\mathcal{V}$. Applying 3D instance segmentation like Mask3D [83] yields instance masks =$\mathcal{M}=\{M_{1},M_{2},\ldots M_{K}\}$, where $K$ denotes the total number of objects in the scene.
For the BEV image $\mathcal{I}_{\bm{b}}$, we begin by projecting the 3D segmentation masks $\mathcal{M}$ onto the $xy$ plane. Then we extract the center coordinates of the resulting bounding box, denoted as $\bm{C}^{\bm{x}\bm{y}}=\{C^{xy}_{1},C^{xy}_{2},\ldots,C^{xy}_{K}\}$. These coordinates are subsequently overlaid onto the BEV image.
For selected egocentric video frames $\mathcal{V}^{*}$, the 3D instance masks $\mathcal{M}$ are projected onto each frame according to its corresponding camera pose. For each frame $i$, we extract the centroid of the 2D mask of each object as its 2D marker, defined as $\bm{C}^{\bm{u}\bm{v}}_{i}=\{C^{uv}_{i,1},C^{uv}_{i,2},\ldots,C^{uv}_{i,K}\}$ , where $C^{uv}_{i,k}$ represents the 2D marker for the $k$-th object in the $i$-th frame. The processed 2D frames with markers and the BEV image with markers are defined as follows:
$\mathcal{V}^{*\prime}$ =$=\left\{\mathcal{F}\left(I_{i},\ \bm{C}_{i}^{\bm{u}\bm{v}}\right)\mid i=s_{1},s_{2},\dots,s_{n}\right\}$ (3)
$\mathcal{I}_{b}^{\prime}$ $=\mathcal{F}\left(\mathcal{I}_{\bm{b}},\ \bm{C}^{\bm{x}\bm{y}}\right)$ (4)
Here, $\mathcal{F}(\cdot)$ represents the operation of overlaying markers onto images, $\mathcal{V}^{*\prime}$ and $\mathcal{I}_{b}^{\prime}$ denote the video frames and BEV image with STO-markers. Notably, the 2D markers in each frame and their corresponding 3D markers in the BEV image maintain spatial alignment, ensuring that they represent the same objects. Moreover, $\bm{C}^{\bm{u}\bm{v}}_{i}$ remains consistent across frames at the object level, preserving temporal coherence.
3.2 Unlocking VLMs with Zero-shot Prompts
We evaluate zero-shot VLMs, focusing on powerful closed-source models (e.g., GPT-4o) to test GPT4Scene’s effectiveness in 3D scene understanding. This process is termed “unlocking”, which enables training-free 3D scene comprehension in VLMs via prompts. Specifically, we input $\mathcal{V}^{*\prime}$ and $\mathcal{I}_{b}^{\prime}$. To reduce the cost, we stitch the images from $\mathcal{V}^{*\prime}$ into a single large image. We evaluate three tasks: (1) In 3D question answering, the goal is to answer scene-related questions like, “What is the color of the floor?”. (2) In dense captioning, the task is to describe a specific object, such as “Describe the object represented by $C_{5}$.” (3) In visual grounding, the aim is to identify the object ID from a description, like “What is the ID of the black chair next to the window?” While question answering is independent of object labels, dense captioning, and visual grounding require object markers. These tasks involve detecting objects and filtering based on the IoU of their bounding boxes. Consistent with Chat-Scene [39] and Robin3D [49], we use Mask3D segmentation results as predicted bounding boxes to calculate the IoU.
In addition to traditional tasks, we have conducted further experiments in this zero-shot setting. Qualitative results are shown in Figure 3. By inputting $\mathcal{V}^{*\prime}$ and $\mathcal{I}_{b}^{\prime}$, VLMs can understand the global features of indoor scenes. At this point, GPT-4o can still accept additional first-person perspective frames, allowing it to understand the current position in the scene to plan the following action. Additionally, using GPT-4o as an agent, VLMs can determine the task type based on the given question and choose the appropriate prompt. Therefore, the GPT4Scene framework shows excellent potential as a core technology for the next generation of embodied intelligence.
3D Question Answering (Scene Level) 3D Dense Caption (Object Level) 3D Visual Grounding (Object Level)
ScanQA SQA3D Scan2cap Multi3DRef ScanRefer
26,138 26,623 35,056 41,408 35,061
Total Samples: 164,286
Table 1: Text Annotations of ScanAlign. We obtained the text annotations of ScanAlign by diversifying the text annotations related to ScanNet, resulting in 165K text annotations.
3.3 Enhancing VLMs with ScanAlign Fine-Tuning
Zero-shot prompts can unlock the 3D understanding capabilities of powerful VLMs, but as shown in  Table 2, this approach does not improve smaller VLMs. Therefore, we aim to enhance open-source, smaller VLMs through fine-tuning. We first construct an indoor scene dataset, ScanAlign, with egocentric, BEV images and text annotations based on the ScanNet [24]. The dataset includes three 3D vision-related tasks represented as $(\mathcal{V}^{*\prime},\mathcal{I}_{b}^{\prime},T)$. The visual input consists of selected video frames with STO-markers and BEV images, and $T$ denotes text annotations derived from five ScanNet annotations [24], as shown in Table 1. We use prompts to randomly vary the annotation format to increase annotation diversity, with further details in the supplementary materials. The dataset contains 165K annotations in total.
Since our method does not require additional modality alignment steps, we can directly perform single-stage instruction fine-tuning on the ScanAlign dataset to enhance the model’s 3D spatial understanding capabilities. During the training phase, the training loss is the Cross-Entropy loss of the language model. The goal is to optimize the learnable parameters, denoted as $\theta$, by minimizing the negative log-likelihood of the target answer, $t^{a}$; we unify the system messages and the user’s question as $t^{q}$. Therefore, the loss function can be expressed as follows:
$\mathcal{L}(\theta)=-\sum_{i=1}^{k}\log P(t^{a}_{i}|t^{a}_{[1,\ldots,i-1]},t^{q}),$ (5)
$k$ represents the number of tokens in the response sequence, and $t^{a}[1,..,i-1]$ denotes the previous $i-1$ tokens in the response. The set of learnable parameters $\theta$ is the vision language projection layers.
After fine-tuning with ScanAlign, we observed significant performance gains for smaller models, which proves that our approach effectively enhances the intrinsic 3D perception capabilities of visual language models (VLMs). We also conducted a series of ablation experiments to validate this finding.
4 Experiments
Zero-shot 3D QA ROUGE@ScanQA EM-1@SQA3D
VID +GPT4Scene VID +GPT4Scene
Open-sourced VLM Based Model
InternVL2-8B [22] 34.3 33.7-$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}-0.6}}}$ 33.0 31.4-$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}-1.6}}}$
MiniCPM-V-2.6-8B [111] 31.5 32.1$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.6}}}$ 42.6 43.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.7}}}$
Qwen2-VL-2B [96] 28.2 28.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+0.2}}}$ 35.7 34.8-$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}-0.9}}}$
Qwen2-VL-7B [96] 29.3 31.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+2.4}}}$ 40.7 41.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+1.0}}}$
[…]
Closed-sourced VLM Based Model
[…]
3D LLM Based Model Pre SOTA Pre SOTA
Chat-Scene [39] 41.6 54.6
Table 2: Zero-shot Setting Results. GPT4Scene yielded limited enhancements for smaller models (2B, 7B, and 8B), while demonstrating significant improvements for larger architectures (72B, closed-source models). This disparity spurred further exploration of fine-tuning strategies for smaller-scale VLMs.
3D Question Answering Point Encoder Vision Encoder ScanQA (val) SQA3D (val)
Methods BLEU-1 BLEU-4 METEOR ROUGE CIDEr EM-1 EM-R1
Task-Specific Model
ScanQA [5] 30.2 10.1 13.1 33.3 64.9 - -
SQA3D [66] - - - - - 46.6 -
3D-VLP [48] 30.5 11.2 13.5 34.5 - - -
3D-Vista [127] - - 13.9 35.7 - 48.5 -
3D LLM Based Model
Chat-3D [100] 29.1 6.4 11.9 28.5 53.2 - -
Chat-3D v2 [38] 38.4 7.3 16.1 40.1 77.1 - -
3D-LLM [37] 39.3 12.0 14.5 35.7 69.4 - -
LL3DA [20] - 13.5 15.9 37.3 76.8 - -
PQ3D [128] - - - - - 47.1 -
LEO [40] - 11.5 16.2 39.3 80.0 50.0 52.4
Chat-Scene [39] 43.2 14.3 18.0 41.6 87.7 54.6 57.5
Vision LLM Based Model
Qwen2-VL-7B [96] 27.8 3.0 11.4 29.3 53.9 40.7 46.7
Qwen2-VL-7B (GPT4Scene) 43.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+15.6}}}$ 14.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+11.6}}}$ 17.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+6.3}}}$ 43.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.2}}}$ 90.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+37.0}}}$ 57.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.7}}}$ 60.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.1}}}$
Qwen2-VL-7B (GPT4Scene)-HD 41.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.1}}}$ 15.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+12.9}}}$ 17.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+6.2}}}$ 43.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+14.2}}}$ 89.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+36.0}}}$ 57.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.5}}}$ 60.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+13.5}}}$
Qwen2-VL-7B (GPT4Scene)-HDM 44.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.6}}}$ 15.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+12.5}}}$ 18.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+7.5}}}$ 46.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+17.1}}}$ 96.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+42.4}}}$ 60.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+19.9}}}$ 63.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+16.4}}}$
Table 3: Evaluation of 3D Question Answering on ScanQA [5] and SQA3D [66] datasets. GPT-4o (GPT4Scene) in the zero-shot setting outperforms most 3D LLM models. Fine-tuned with GPT4Scene, Qwen2-VL achieves state-of-the-art performance. The base setting uses $N=8$ frames at $128\times 123$, ”HD” increases resolution to $512\times 490$, and ”HDM” combines this resolution with $N=32$ frames.
In this section, we primarily present the experimental results. First, we evaluate the improvements of GPT4Scene under zero-shot settings in Section 4.1. Following this, we outline the training details in Section 4.2, and subsequently introduce the fine-tuning results in Section 4.3. Finally, Section 4.4 details the ablation study, demonstrating the effectiveness of individual components.
4.1 Zero-shot Setting Results
Can VLMs understand the 3D world directly?
First, we examine GPT4Scene’s zero-shot performance, with the results detailed in Table 2. We begin by comparing a range of open-source models, including InternVL2-8B [22], MiniCPM-V-2.6-8B [111], and the Qwen2-VL family (2B, 7B, and 72B) [96]. Our analysis reveals that for smaller models (2B, 7B, and 8B), the integration of GPT4Scene demonstrates limited effectiveness in enhancing 3D spatial understanding through VLMs. This suggests that the benefits of GPT4Scene are more pronounced in larger models, highlighting the need for further optimization and architectural advancements to bridge this gap.
However, significant improvements are observed with larger models like Qwen2-VL-72B. When testing API-based open-source models Gemini-1.5-Pro [88] and GPT-4o [72], we find that incorporating BEV and STO-markers yields substantial gains, achieving performance comparable to previous state-of-the-art methods Chat-Scene [39]. These results suggest that smaller VLMs face inherent limitations in visual comprehension capabilities, as they cannot directly enhance 3D scene understanding through zero-shot prompting alone. Consequently, we propose constructing a specialized dataset, ScanAlign, and conducting fine-tuning on Qwen2-VL to achieve performance improvements.
4.2 Implementation Details
Our 3D scene understanding benchmark is built upon the ScanNet dataset (1,513 scenes, following its original data splits) and encompasses three core tasks: 3D question answering (evaluated via ScanQA [5] and SQA3D [66]), 3D dense captioning (Scan2Cap [21]), and 3D visual grounding (ScanRefer [12] and Multi3DRef [121]). The data processing pipeline includes reconstructing point clouds with BundleFusion [25] (the original ScanNet method), performing 3D instance segmentation with Mask3D [83] (consistent with Chat-Scene [39] and Robin3D [49]), projecting object centers to bird’s-eye-view coordinates, and mapping 3D masks onto 2D images for marker positioning. 2D marker placement is necessary for dense captioning and visual grounding tasks.
For the experienment config, we sample N=8 frames per video (128×123 resolution). Closed-source visual language models (VLMs) like GPT-4o process stitched 8-frame panoramas, while open-source models (e.g., Qwen2-VL-7B) directly take in sampled frames. The open-source models also support enhanced configurations: HD mode (high resolution, 512×490) and HDM mode (high resolution with multiple frames, 512×490, 32 frames). Training is conducted for one epoch using a base learning rate of 5e-6 with cosine annealing scheduling, completing in approximately 6 hours on 8×A100 GPUs. 2D marker localization is preserved to meet the specific needs of dense captioning and visual grounding tasks.
4.3 Fine-tuning Main Results
We present our experimental results using the ScanAlign dataset for VLM fine-tuning. The results demonstrate that using only Qwen2-VL-7B [96], we achieve state-of-the-art performance across all indoor scene understanding tasks on ScanNet. This validates the effectiveness of GPT4Scene in enhancing VLMs’ capacity for 3D scene comprehension.
3D Question Answering. The 3D question answering results are presented in Table 3. We categorize the evaluated approaches into three groups: (1) classic task-specific models optimized for 3D QA tasks, (2) 3D point cloud LLM-based architectures, and (3) vision-language multimodal LLM systems. Our experimental analysis demonstrates that the baseline Qwen2-VL-7B model without fine-tuning shows constrained capability in 3D QA scenarios. Through systematic fine-tuning with GPT4Scene, the enhanced Qwen2-VL-7B(GPT4Scene) achieves state-of-the-art performance in 3D visual question answering, surpassing all comparative methods across evaluation metrics. Quantitative comparisons reveal 56.1% improvement in BLEU-1 (27.8 → 43.4) and 68.6% enhancement in CIDEr (53.9 → 90.9) on the ScanQA benchmark compared to the original Qwen2-VL-7B implementation. For SQA3D evaluations, we observe a 41.0% increase in EM-1 scores (40.7 → 57.4). Furthermore, under HD (High-Resolution) and HDM (High-Resolution & multi frames) configurations, the model exhibits additional performance gains.
3D Dense Caption IoU@0.25 IoU@0.5
Methods BLEU-4 ROUGE BLEU-4 ROUGE
Task-Specific Model
Scan2Cap [21] 34.2 55.3 23.3 44.5
3DJCG [8] 40.2 59.2 31.0 50.8
X-Trans2Cap [114] 35.7 54.7 25.1 45.3
3D-VisTA [127] 36.5 57.6 34.0 54.3
Vote2Cap-DETR [19] 39.3 59.3 34.5 54.4
3D LLM Based Model
LL3DA [20] 41.4 59.5 36.8 55.1
LEO [40] 36.9 57.8
Chat-Scene [39] 38.2 60.6 36.3 58.1
Robin3D [49] 38.4
Vision LLM Based Model
Qwen2-VL-7B [96] 3.8 24.7 3.8 24.6
Qwen2-VL-7B (GPT4Scene) 36.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.5}}}$ 57.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.9}}}$ 34.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+30.4}}}$ 55.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+30.6}}}$
Qwen2-VL-7B (GPT4Scene)-HD 40.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+36.6}}}$ 60.2$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+35.5}}}$ 37.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+34.1}}}$ 57.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+33.1}}}$
Qwen2-VL-7B (GPT4Scene)-HDM 43.1$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+39.3}}}$ 61.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+37.2}}}$ 40.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+36.8}}}$ 59.3$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+34.7}}}$
Table 4: Evaluation of 3D Dense Caption on Scan2Cap [21]. Our results outperform those of existing 3D LLM based models.
Figure 3: Qualitive Results. We demonstrate the qualitative performance of GPT-4o under zero-shot settings using GPT4Scene prompts, highlighting two distinctive application cases that exemplify the versatility of GPT4Scene.
3D Dense Caption & Visual Grounding. We further assess the model’s capabilities in dense captioning and visual grounding tasks, which differ fundamentally from question answering by requiring explicit grounding mechanisms between textual descriptions and 3D spatial structures. As evidenced by the experimental results in Table 4 and Table 6, our GPT4Scene-finetuned Qwen2-VL-7B demonstrates substantial gains in both 3D caption generation and object localization accuracy compared to baseline vision-language models (VLMs). The performance improvements become particularly pronounced under HD and HDM configurations, where our approach establishes new state-of-the-art results across all evaluation metrics, exceeding existing methods by significant margins.
3D Visual Grounding ScanRefer Multi3DRef
Methods Acc@0.25 Acc@0.50 all F1@0.25 all F1@0.50
Task-Specific Model
3DVG-Transformer [122] 47.6 34.7 25.5
3DJCG [8] 49.6 37.3 26.6
D3Net [13] 37.9 32.2
M3DRef-CLIP [121] 51.9 44.7 42.8 38.4
3D LLM Based Model
Chat-Scene [39] 55.5 50.2 57.1 52.4
Vision LLM Based Model
Qwen2-VL-7B [96] 5.4 5.1 21.1 19.9
Qwen2-VL-7B (GPT4Scene) 40.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+35.1}}}$ 36.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+31.6}}}$ 45.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+24.3}}}$ 42.1+22.$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+22.2}}}$
Qwen2-VL-7B (GPT4Scene)-HD 50.9$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+45.5}}}$ 46.4$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+41.3}}}$ 53.7$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.6}}}$ 50.0$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+30.1}}}$
Qwen2-VL-7B (GPT4Scene)-HDM 62.6$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+51.9}}}$ 57.0$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+32.3}}}$ 64.5$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+43.4}}}$ 59.8$\,{}_{\text{{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}+39.9}}}$
Table 5: Evaluation of 3D Visual Grounding on ScanRefer [12] and Multi3DRef [121]. Our method reaches SOTA performance over all methods.
3D QA GPT Score Combat Situation GPT Score
Methods Win Tie Lose Score
VLM vs 3D LLM
Qwen2-VL-7B [96] vs Chat-Scene [39] 74 243 683 465
Qwen2-VL-7B(GPT4Scene) vs Chat-Scene 543 145 312 1774
Table 6: GPT Score evaluation on 3D question answering. We evaluated the GPT Score on ScanQA [5], comparing our model’s outputs with Chat-Scene [39]. The fine-tuned model achieved significantly higher GPT Scores than current SOTA methods.
Qualitative Results. We present qualitative results in Figure 3, conducted in a zero-shot setting on GPT-4o. Stitched frames provide an overview of the scene, while individual frames capture details and actions. Beyond standard tasks like object captioning, spatial description, and counting, GPT4Scene handles embodied tasks, such as directing users to retrieve paper from a nearby bookshelf. In the last row, pink-highlighted segments indicate BEV images as input, enhancing navigation capabilities. GPT4Scene also excels in navigation and patrolling tasks, performing industrial inspections by observing machine indicators.
GPT Score: New Evaluation Way. We propose a novel GPT Score for 3D QA assessment, comparing two methods with the state-of-the-art 3D LLM Chat-Scene [39] as baseline. In the comparison with Qwen2-VL-7B [96], we input both models’ outputs and the ground truth into GPT-4o [72] to evaluate which response is better. We select 1,000 questions from ScanQA and define ”win,” ”tie,” and ”lose” to indicate whether Qwen2-VL-7B’s answer is superior to, equivalent to, or inferior to Chat-Scene’s [39]. These outcomes are assigned scores of 3, 1, and -1 points respectively (analogous to football scoring), which are aggregated to derive a comprehensive GPT Score. Our results show that Qwen2-VL-7B (without fine-tuning) underperforms Chat-Scene, while its fine-tuned version (GPT4Scene) significantly outperforms it, demonstrating its effectiveness.
Small Objects Sets ScanQA
ROUGE
Chat-Scene 37.5
GPT-4o (GPT4Scene) 35.4
Qwen2-VL-7B (GPT4Scene) 39.4
(a) Ablation on Small Objects.
Precision and Size of STO-markers ScanQA
ROUGE
Qwen2-VL-7B (GPT4Scene) 43.6
- Delete 30% STO-markers 42.7
- Adaptive size adjustment 43.0
(b) Ablation on STO-markers.
BEV Reconstruction Quality ScanQA
ROUGE
Qwen2-VL-7B (GPT4Scene) 43.6
- SLAM3R [63], 50-frame intervals 42.4
- SLAM3R [63], 100-frame intervals 41.9
- SLAM3R [63], 200-frame intervals 43.2
(c) Ablation on BEV Reconstruction Quality.
Table 7: Overall Performance Comparison. We investigated GPT4Scene’s performance on small objects and its robustness in handling STO-markers and BEV reconstruction quality, demonstrating that our model exhibits strong performance and robustness.
Ablation ScanQA Multi3DRef@0.5
ROUGE CIDEr MT ALL
Qwen2-VL-7B (GPT4Scene) 43.6 90.9 36.3 42.1
w/o BEV Image 42.3 87.1 27.8 32.1
w/o STO markers 42.8 88.4 - -
w/o BEV Image & STO markers 41.7 85.0 - -
Table 8: Ablation Study on our modules. Removing BEV images or STO-markers leads to a performance decrease, and further removal of STO-markers causes an additional decrease.
4.4 Ablation Study
In this section, we conduct ablation studies to validate the effectiveness of GPT4Scene. First, we evaluate its robustness, including performance on small objects, followed by analyzing the robustness of STO-markers and reconstruction quality. Next, we perform module-wise ablation to assess individual components. Finally, we investigate the impact of frame intervals and resolution settings.
Ablation on Robustness Analysis. We tested the performance of GPT4Scene on small objects by selecting 1,000 instances, such as cups and towels. The results demonstrate that the fine-tuned Qwen2-VL-7B [96] model with GPT4Scene exhibits significantly improved performance. Following this, we analyzed the robustness of STO-markers. Two modifications were implemented: first, removing 30% of the markers, and second, dynamically adjusting marker sizes based on object dimensions. Experimental results indicate that both operations had minimal impact on overall performance, confirming the robustness of STO-markers. The third experiment investigated the impact of BEV (Bird’s Eye View) map reconstruction quality on results. Originally, we used BundleFusion [25] from ScanNet [24] for reconstruction. We replaced this with SLAM3R [63], a real-time reconstruction method, and tested settings with frame intervals of 50, 100, and 200. The results show that while reconstruction quality slightly affects performance, the influence is negligible. This suggests that the reconstructed BEV primarily serves to provide global scene context rather than precise geometric details.
Ablation on Our Modules. Table 2 demonstrates that BEV images and STO-markers enhance spatial understanding under zero-shot settings. We further validate this observation in the fine-tuning scenario through evaluations on both 3D question answering and 3D visual grounding tasks. It should be noted that STO-markers are essential for visual grounding——when they are removed, validation can only be performed on the question answering task. As shown in Table 8, both BEV images and STO-markers contribute significantly to the performance improvements.
Ablation on Frames and Resolution. Here, we conduct ablation studies on additional factors using 3D question answering and visual grounding as benchmark tasks. The results are shown in Table 9. We define three configurations: the base mode with 8 frames and longer side resolution of 128, the HD (high-resolution) mode with 8 frames at 512 resolution, and the HDM (high-resolution multi-frame) mode with 32 frames at 512 resolution. The first three rows demonstrate that image resolution substantially impacts visual grounding performance while providing limited improvement for the QA task. The final three rows reveal that frame quantity augmentation enhances indoor scene comprehension, showing more pronounced gains in grounding performance compared to marginal QA improvements.
Num Resolution ScanQA ScanRefer
Frames ROUGE CIDEr Acc0.25 Acc0.5
8 128 (base) 43.6 90.9 40.5 36.7
256 43.8 90.0 49.2 44.8
512 (HD) 43.6 89.9 50.9 46.4
16 512 45.4 93.4 58.6 53.4
32 512 (HDM) 46.5 96.3 62.6 57.0
Table 9: Ablation Study on frames and resolution. Results show that the number of frames affects both QA and grounding, while resolution significantly impacts grounding.
5 Conclusion
We propose GPT4Scene, a framework enhancing Vision-Language Models (VLMs) to understand 3D scenes directly from visual inputs. By reconstructing 3D point clouds for Bird’s Eye View (BEV) images and aligning video frames with spatial-temporal object (STO) markers, we enable global-local scene comprehension. GPT4Scene achieves state-of-the-art 3D QA performance with zero-shot GPT-4o and fine-tuned smaller VLMs (e.g., Qwen2-VL) using our ScanAlign dataset. Fine-tuned models even excel with raw video inputs, proving effective 3D understanding.
Appendix A Prompts of Closed-source VLMs
Here, we present the prompts used by GPT4Scene in the closed-source VLMs (GPT-4o), as illustrated in Figure 4. The process begins with a general system prompt, which outlines the overview of two images provided as input. The first image is a stitched 2D view captured from a video, with dimensions of $2\times 4$. The second image represents a BEV (Bird’s Eye View). Subsequently, we perform evaluations across various tasks and benchmarks, with each benchmark associated with a specific prompt. We take the 3D question-answering task on ScanQA as an example. The benchmark prompt consists of three parts:
  1. 1.
    Important Guidelines: It clarifies that while we provide object IDs for reference, they cannot be directly used when answering questions. Additionally, it specifies adapting the response style to match that of ScanQA. Since ScanQA’s responses are typically short single words, we aim to keep the answers concise within the benchmark prompt, targeting 1-5 words.
  2. 2.
    Answer Format: In this part, we use a standardized regularized format to structure the answers, which helps improve accuracy when addressing questions.
  3. 3.
    Examples: we include two example cases.
Our zero-shot prompting process is illustrated in the bottom-left corner Figure 4. The system prompt and ScanQA prompt are used as the system message. In the user message, we input the stitched image and BEV image. Finally, the query message includes the question. The responses generated through this process require further refinement, as depicted in the bottom-right corner of Figure 4. First, we remove the regularized formatting from the answers. Next, we clean the answers by addressing singular/plural forms and case sensitivity. This final step ensures that we obtain the refined answers.
Task type Model
Qwen2-VL Ours
Action Sequence 85.5 82.0
Action Prediction 69.5 70.5
Action Antonym 83.0 86.0
Fine-grained Action 51.5 51.5
Unexpected Action 82.0 78.0
Object Existence 87.5 88.5
Object Interaction 82.0 81.5
Object Shuffle 41.0 45.0
Moving Direction 42.0 40.0
Action Localization 65.0 66.5
Scene Transition 93.5 94.0
Action Count 47.5 43.5
Moving Count 69.5 71.5
Moving Attribute 90.0 88.5
State Change 48.0 49.0
Fine-grained Pose 63.0 63.5
Character Order 74.5 71.0
Egocentric Navigation 39.5 41.5
Episodic Reasoning 47.0 47.0
Counterfactual Inference 62.5 65.5
Avg 66.2 66.225
Table 10: The result of MVBench [57]. After fine-tuning with ScanAlign in GPT4Scene, our model shows improved 2D understanding, particularly in object and action metrics.
      Benchmark       Model
      Qwen2-VL       Ours
      MMBench-ENval [115]       82.4       81.2
      MMBench-CNval [115]       81.7       79.9
      MMStar [16]       60.7       57.6
      RealWorldQA [104]       70.1       68.5
      Video-MME [30]       59.8       58.4
Table 11: The result of 2D Multi-modal Benchmark. After fine-tuning with ScanAlign in GPT4Scene, our model’s 2D understanding capabilities did not decline.
Methods BLEU-n Metrics Language Generation Metrics
EM-1 BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE METEOR CIDEr
Task-Specific Model
ScanQA [5] 21.1 30.2 20.4 15.1 10.1 33.3 13.1 64.9
3D-VLP [48] 21.7 30.5 21.3 16.7 11.2 34.5 13.5 67.0
3D-Vista [127] 13.9 35.7
3D LLM Based Model
3D-LLM [37] 20.5 39.3 25.2 18.4 12.0 35.7 14.5 69.4
LL3DA [20] 13.5 37.3 15.9 76.8
LEO [40] 24.5 11.5 39.3 16.2 80.0
Scene-LLM [31] 27.2 43.6 26.8 19.1 12.0 40.0 16.6 80.0
Chat-Scene [39] 21.6 43.2 29.1 20.6 14.3 41.6 18.0 87.7
Vision LLM Based Model
Qwen2-VL-7B [96] 19.0 27.8 13.6 6.3 3.0 29.3 11.4 53.9
Qwen2-VL-7B (GPT4Scene) 25.5 43.4 29.3 20.9 14.6 43.6 17.7 90.9
Qwen2-VL-7B (GPT4Scene)-HD 26.3 41.9 28.6 21.6 15.9 43.6 17.6 89.9
Qwen2-VL-7B (GPT4Scene)-HDM 28.2 44.4 30.3 22.3 15.5 46.5 18.9 96.3
Table 12: Full Evaluation of 3D Question Answering on ScanQA [5].
      Methods       Test Set       Avg.(EM-1)       EM-R1
      What       Is       How       Can       Which       Others
      Task-Specific Model
      ClipBERT [66]       30.2       60.1       38.7       63.3       42.5       42.7       43.3       –
      SQA3D [66]       31.6       63.8       46.0       69.5       43.9       45.3       46.6       –
      3D-VisTA [127]       34.8       63.3       45.4       69.8       47.2       48.1       48.5       –
      3D LLM Based Model
      PQ3D [128]       37.1       61.3       44.5       60.9       47.0       45.1       47.1       –
      LEO [40]       –       –       –       –       –       –       50.0       52.4
      Scene-LLM [31]       40.9       69.1       45.0       70.8       47.2       52.3       54.2       –
      Chat-Scene [39]       45.4       67.0       52.0       69.5       49.9       55.0       54.6       57.5
      Vision LLM Based Model
      Qwen2-VL-7B [96]       29.0       59.2       33.4       50.5       44.2       43.2       40.7       46.7
      Qwen2-VL-7B (GPT4Scene)       50.7       70.9       48.0       70.5       52.9       59.3       57.4       60.7
      Qwen2-VL-7B (GPT4Scene)-HD       51.4       69.1       50.2       69.4       51.3       57.9       57.2       60.4
      Qwen2-VL-7B (GPT4Scene)-HDM       55.9       69.9       50.8       68.7       53.3       60.4       59.4       62.4
Table 13: Full Evaluation of 3D Question Answering on SQA3D [66].
Appendix B 2D Multi-modal Benchmark
We tested the fine-tuned Qwen2-VL-7B (GPT4Scene) model on 2D image and video multimodal large models. Table 10 shows the results of MVBench. As we can see, our model shows improvement for the object and action metrics, indicating that the model fine-tuned with ScanAlign is better at handling spatial variations and the information of objects in the scene. Table 11 presents the results on other benchmarks, where we can observe that after our fine-tuning, the model’s ability to understand images and videos did not decline significantly. It demonstrates the effectiveness of ScanAlign.
Appendix C Qualitative Results
Figures 5, 6, 7 and 8 present our qualitative results obtained from Qwen2-VL after fine-tuning on ScanAlign. Figures 5 and 6 demonstrate 3D question answering performance using only unannotated video inputs as context. Our observations indicate the model’s capability to effectively address queries from both ScanQA [5] and SQA3D [66] benchmarks. Meanwhile, Figures 7 and 8 exhibit the model’s competence in 3D dense captioning and 3D visual grounding tasks requiring annotated inputs. These outcomes substantiate that our framework can generate precise responses through visual information processing alone, without dependency on 3D point cloud inputs. Here we emphasize that post-training with the GPT4Scene framework, the model achieves accurate QA performance using only pure video input, thereby demonstrating GPT4Scene’s effectiveness in enhancing visual comprehension capabilities.
Appendix D Full Quantitive Results
Here, we present the complete metrics for all five benchmarks. Table 6(c) and Table 13 show results for ScanQA [5] and SQA3D [66]. Table 14 provides the full results for Scan2Cap [21], while Table 15 and Table 16 present the results for ScanRefer [12] and Multi3DRef [121]. Our model significantly improves across all benchmarks, highlighting that only pure vision language models can understand 3D scenes effectively.
Methods IoU@0.25 IoU@0.5
CIDEr BLEU-4 METEOR ROUGE CIDEr BLEU-4 METEOR ROUGE
Task-Specific Model
Scan2Cap [21] 56.8 34.2 26.3 55.3 39.1 23.3 22.0 44.5
3DJCG [8] 64.7 40.2 27.7 59.2 49.5 31.0 24.2 50.8
X-Trans2Cap [114] 61.8 35.7 26.6 54.7 43.9 25.1 22.5 45.3
D3Net [13] 62.6 35.7 25.7 53.9
3D-VLP [48] 70.7 41.0 28.1 59.7 54.9 32.3 24.8 51.5
Vote2Cap-DETR [19] 71.5 39.3 28.3 59.3 62.6 35.7 25.7 53.9
3D-VisTA [127] 71.0 36.5 28.4 57.6 66.9 34.0 27.1 54.3
3D LLM Based Model
LL3DA [20] 74.2 41.4 27.8 59.5 65.2 36.8 26.0 55.1
LEO [40] 68.4 36.9 27.7 57.8
Chat-Scene [39] 81.9 38.2 29.0 60.6 77.2 36.3 28.0 58.1
Robin3D [49] 87.2 38.4
Vision LLM Based Model
Qwen2-VL-7B [96] 0.0 3.8 16.7 24.7 0.0 3.8 16.5 24.6
Qwen2-VL-7B (GPT4Scene) 63.8 36.3 26.5 57.6 60.6 34.2 25.6 55.2
Qwen2-VL-7B (GPT4Scene)-HD 79.1 40.4 28.3 60.2 74.4 37.9 27.3 57.7
Qwen2-VL-7B (GPT4Scene)-HDM 91.7 43.1 29.3 61.9 86.3 40.6 28.2 59.3
Table 14: Full Evaluation of 3D Dense Caption on Scan2Cap [21].
Methods Unique Multiple Overall
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
Task-Specific Model
ScanRefer [12] 76.3 53.5 32.7 21.1 41.2 27.4
TGNN [41] 68.6 56.8 29.8 23.2 37.4 29.7
SAT [114] 73.2 50.8 37.6 25.2 44.5 30.1
InstanceRefer [113] 75.7 64.7 29.4 23.0 38.4 31.1
3DVG-Transformer [122] 81.9 60.6 39.3 28.4 47.6 34.7
MVT [42] 77.7 66.4 31.9 25.3 40.8 33.3
3D-SPS [64] 84.1 66.7 40.3 29.8 48.8 37.0
ViL3DRel [18] 81.6 68.6 40.3 30.7 47.9 37.7
3DJCG [8] 83.5 64.3 41.4 30.8 49.6 37.3
D3Net [13] 72.0 30.1 37.9
BUTD-DETR [45] 84.2 66.3 46.6 35.1 52.2 39.8
HAM [14] 79.2 67.9 41.5 34.0 48.8 40.6
3DRP-Net [99] 83.1 67.7 42.1 32.0 50.1 38.9
3D-VLP [48] 84.2 64.6 43.5 33.4 51.4 39.5
EDA [103] 85.8 68.6 49.1 37.6 54.6 42.3
M3DRef-CLIP [121] 85.3 77.2 43.8 36.8 51.9 44.7
3D-VisTA [127] 81.6 75.1 43.7 39.1 50.6 45.8
ConcreteNet [93] 86.4 82.1 42.4 38.4 50.6 46.5
DOrA [101] 52.8 44.8
3D LLM Based Model
Chat-Scene [39] 89.6 82.5 47.8 42.9 55.5 50.2
Robin3D [49] 60.8 55.1
Vision LLM Based Model
Qwen2-VL-7B [96] 6.3 6.3 5.1 4.8 5.4 5.1
Qwen2-VL-7B (GPT4Scene) 65.5 61.2 34.8 31.1 40.5 36.7
Qwen2-VL-7B (GPT4Scene)-HD 77.5 71.9 44.9 40.6 50.9 46.4
Qwen2-VL-7B (GPT4Scene)-HDM 90.3 83.7 56.4 50.9 62.6 57.0
Table 15: Full Evaluation of 3D Visual Grounding on ScanRefer [12].
Methods ZT w/o D ZT w/ D ST w/o D ST w/ D MT ALL
F1 F1 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5 F1@0.25 F1@0.5
Task-Specific Model
3DVG-Trans+ [122] 87.1 45.8 27.5 16.7 26.5 25.5
D3Net (Grounding) [13] 81.6 32.5 38.6 23.3 35.0 32.2
3DJCG (Grounding) [8] 94.1 66.9 26.0 16.7 26.2 26.6
M3DRef-CLIP [121] 81.8 39.4 53.5 47.8 34.6 30.6 43.6 37.9 42.8 38.4
3D LLM Based Model
Chat-Scene [39] 90.3 62.6 82.9 75.9 49.1 44.5 45.7 41.1 57.1 52.4
Vision LLM Based Model
Qwen2-VL-7B [96] 84.8 58.5 20.1 19.1 14.7 13.5 16.8 15.4 21.1 19.9
Qwen2-VL-7B (GPT4Scene) 85.2 61.4 60.1 55.1 37.7 34.4 39.4 36.3 45.4 42.1
Qwen2-VL-7B (GPT4Scene)-HD 93.6 81.8 72.5 66.2 46.6 42.9 41.8 38.9 53.7 50.0
Qwen2-VL-7B (GPT4Scene)-HDM 97.4 84.4 85.0 77.7 59.9 55.1 48.6 44.6 64.5 59.8
Table 16: Full Evaluation of 3D Visual Grounding on Multi3DRef [121].
Figure 4: Prompts of Closed-sourse VLMs. We show the prompts used for GPT-4o (GPT4Scene), which consist of a system prompt and a benchmark prompt. After generating responses, we further refine them.
Figure 5: Qualitive Results: Question Answering. We provide videos without object annotations.
Figure 6: Qualitive Results: Question Answering. We provide videos without object annotations.
Figure 7: Qualitive Results: 3D Dense Caption and Visual Grounding. We provide videos with object annotations.
Figure 8: Qualitive Results: 3D Dense Caption and Visual Grounding. We provide videos with object annotations.